find the magnitude of vectorF =2i+3j+5k and also calculate the angles of vector makes with x-axis, y-axis and z-axis
1
Expert's answer
2019-10-28T11:00:47-0400

The magnitude of vector is equal to

"a=\sqrt{4+9+25}=6.2"


the angles of vector makes with x-axis

"\cos{\frac{2}{6.2}}=0.3226"

Angle with x-axis is equal to 70°


the angles of vector makes with y-axis

"\cos{\frac{3}{6.2}}=0.4823"

Angle with y-axis is equal to 60°


the angles of vector makes with z-axis

"\cos{\frac{5}{6.2}}=0.8064"

Angle with z-axis is equal to 35°